# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v4

@@ C.1 Proof of Theorem 5
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ $=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}$ (70)
$=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\operatorname{vec}\Biggl(\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]\Biggr)}_{2}^{2}$ (71)
$=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\operatorname{vec}\Biggl(\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr)}_{2}^{2}$ (72)
$=\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{\top}\otimes\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\Biggr]\bm{u}_{i}}_{2}^{2}.$ (73)
@@
$\bm{K}$ $=\begin{bmatrix}\bm{I}\otimes\prod_{j=2}^{L}\bm{W}^{*}_{j}\big|{\bm{W}_{1}^{*}}^{\top}\otimes\Bigl(\prod_{j=3}^{L}\bm{W}^{*}_{j}\Bigr)\big|&\dots&\big|\Bigl(\prod_{j=1}^{L-1}\bm{W}_{j}^{*}\Bigr)^{\top}\otimes\bm{I}\\ \end{bmatrix}$, (74)
$\bm{u}$ $=[\bm{u}_{1}^{\top}\quad\bm{u}_{2}^{\top}\quad\cdots\quad\bm{u}_{L}^{\top}]^{\top}.$ (75)
Then,
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ $=\max_{\bm{u}:\norm{\bm{u}}_{2}=1}2\norm{\bm{K}\bm{u}}_{2}^{2}$ (76)
$=\sigma_{\max}(\bm{K}^{\top}\bm{K}).$ (77)
Note that $\sigma_{\max}(\bm{K}^{\top}\bm{K})=\sigma_{\max}(\bm{K}\bm{K}^{\top})$, and for any two block matrices $\bm{A}$ and $\bm{B}$ such that
$\bm{A}=\begin{bmatrix}\bm{A}_{1}\quad\bm{A}_{2}\quad\dots\quad\bm{A}_{L}\end{bmatrix}\in\mathbb{R}^{M_{1}\times d},\quad\bm{B}=\begin{bmatrix}\bm{B}_{1}\\ \vdots\\ \bm{B}_{L}\\ \end{bmatrix}\in\mathbb{R}^{d\times M_{2}}$ (78)
$\bm{AB}=\sum_{i=1}^{L}\bm{A}_{i}\bm{B}_{i},\quad\bm{AB}\in\mathbb{R}^{M_{1}\times M_{2}}.$ (79)
Furthermore, for any matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ such that the matrix products $\bm{AB}$ and $\bm{CD}$ are well defined, we have
$(\bm{A}\otimes\bm{C})(\bm{B}\otimes\bm{D})=\bm{AB}\otimes\bm{CD}.$ (80)
Using the fact that $(\bm{A}\otimes\bm{C})^{\top}=\bm{A}^{\top}\otimes\bm{C}^{\top}$ together with the previous property, it follows that
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)=2\sigma_{\max}\Biggl(\sum_{i=1}^{L}\bm{B}_{i}^{\top}\bm{B}_{i}\otimes\bm{A}_{i}\bm{A}_{i}^{\top}\Biggr),$ (81)
where $\bm{A}_{k}=\prod_{i=k+1}^{L}\bm{W}_{i}^{*}$ and $\bm{B}_{k}=\prod_{i=1}^{k-1}\bm{W}_{i}^{*}$.
@@ C.2 Proof of Corollary 25
According to (14), for any $(\bm{L}^{*},\bm{R}^{*})\in\Omega$,
$\nabla^{2}\mathcal{L}(\bm{L}^{*},\bm{R}^{*})[\bm{U},\bm{V}]=2\norm{\bm{L}^{*}\bm{U}^{\top}+\bm{V}\bm{R}^{*\top}}_{F}^{2}.$ (82)
Then,
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{L}^{*},\bm{R}^{*}))$ $=\max_{\begin{subarray}{c}\bm{U},\bm{V}\\ \|\bm{U}\|_{F}^{2}+\|\bm{V}\|_{F}^{2}=1\end{subarray}}2\norm{\bm{L}^{*}\bm{U}^{\top}+\bm{V}\bm{R}^{*\top}}_{F}^{2}$ (83)
$\leq\max_{\begin{subarray}{c}\bm{U},\bm{V}\\ \|\bm{U}\|_{F}^{2}+\|\bm{V}\|_{F}^{2}=1\end{subarray}}2\bigl(\norm{\bm{L}^{*}\bm{U}^{\top}}_{F}+\norm{\bm{V}\bm{R}^{*\top}}_{F}\bigr)^{2}$ (84)
$=\max_{\begin{subarray}{c}\bm{u},\bm{v}\\ \|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}=1\end{subarray}}2\Bigl(\|(\bm{I}\otimes\bm{L}^{*})\bm{u}\|_{2}+\|(\bm{R}^{*}\otimes\bm{I})\bm{v}\|_{2}\Bigr)^{2}$ (85)
$\leq\max_{\begin{subarray}{c}\bm{u},\bm{v}\\ \|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}=1\end{subarray}}2\Bigl(\sigma_{\max}(\bm{I}\otimes\bm{L}^{*})\|\bm{u}\|_{2}+\sigma_{\max}(\bm{R}^{*}\otimes\bm{I})\|\bm{v}\|_{2}\Bigr)^{2}.$ (86)
We can upper bound the right-hand side of (83) using the triangle inequality. By applying the vectorization trick of the Kronecker product again, we can rewrite (84). Then, noting that for any matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and vector $\bm{x}\in\mathbb{R}^{n}$, $\norm{\bm{Ax}}_{2}\leq\sigma_{\max}(\bm{A})\norm{\bm{x}}_{2}$, we can upper bound the right-hand side of (85). Note that for any matrix $\bm{A}$ and $\bm{B}$, $\sigma_{\max}(\bm{A}\otimes\bm{B})=\sigma_{\max}(\bm{A})\sigma_{\max}(\bm{B})$. Hence,
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{L}^{*},\bm{R}^{*}))\leq\max_{\begin{subarray}{c}\bm{u},\bm{v}\\ \|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}=1\end{subarray}}2\bigl(\sigma_{\max}{(\bm{L}^{*})}\norm{\bm{u}}_{2}+\sigma_{\max}(\bm{R}^{*})\norm{\bm{v}}_{2}\bigr)^{2}.$ (87)
@@
$\lambda_{\max}\!\bigl(\nabla^{2}\mathcal{L}(\bm{L}^{*},\bm{R}^{*})\bigr)\;\leq\;2\bigl(\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}\bigr).$ (88)
Now, we will show that this upper bound is achievable. Let us decompose $\bm{L}^{*}$ as $\bm{U}_{L}\bm{\Sigma}_{L}\bm{V}_{L}^{\top}$ by SVD, and denote by $\bm{u}_{L}$ and $\bm{v}_{L}$ the left and right singular vectors corresponding to the largest singular value, respectively. Moreover, decompose $\bm{R}^{*}$ as $\bm{U}_{R}\bm{\Sigma}_{R}\bm{V}_{R}^{\top}$ by SVD, and denote by $\bm{u}_{R}$ and $\bm{v}_{R}$ the left and right singular vectors corresponding to the largest singular value, respectively. We determine a particular $(\bm{U}^{*},\bm{V}^{*})$ such that it achieves the upper bound while satisfying the constraint $\norm{\bm{U}^{*}}_{F}^{2}+\norm{\bm{V}^{*}}_{F}^{2}=1$. Choose
$\bm{U}^{*\top}=\frac{\sigma_{\max}(\bm{L}^{*})}{\sqrt{\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}}}\bm{v}_{L}\bm{u}_{R}^{\top},\quad\bm{V}^{*}=\frac{\sigma_{\max}(\bm{R}^{*})}{\sqrt{\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}}}\bm{u}_{L}\bm{v}_{R}^{\top}.$ (89)
Using the fact that, for any vectors $\bm{x}$ and $\bm{y}$ $\norm{\bm{x}\bm{y}^{\top}}_{F}^{2}=\norm{\bm{x}}_{2}^{2}\norm{\bm{y}}_{2}^{2}$,
$2\norm{\frac{(\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2})}{\sqrt{\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}}}\bm{u}_{L}\bm{u}_{R}^{\top}}_{F}^{2}=2(\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}).$ (90)
@@ D.1 Visualization of the Contour Map of the Loss Landscape
To study the dynamics of deep matrix factorization, we analyze the trajectories of GD. Previous works have visualized neural network loss landscapes to explore their highly non-convex and non-Euclidean structure (Goodfellow et al., 2014; Li et al., 2018). However, the high-dimensionality prevents full visualization. As a result, only 1-D (line) or 2-D (surface) visualizations are available. In this paper, we focus on contour maps of the loss landscape in the vicinity of a global minimum and a methodology employed in prior studies to generate them.
Contour Plots with Random Projections. We want to visualize the loss landscape around a global minimum $\bm{w}^{*}\in\mathbb{R}^{N}$. We select two random vectors, $\bm{\zeta}$ and $\bm{\gamma}$, from $\mathbb{R}^{N}$. Then, for any $K\subset\mathbb{R}^{2}$, we can define the function $p:K\rightarrow\mathbb{R}:$
$p(x,y)=\mathcal{L}(\bm{w}^{*}+x\bm{\zeta}+y\bm{\gamma}),\quad\forall(x,y)\in K,$ (91)
@@
Scale Invariance and Manifolds. Note that our loss function is scale-invariant, which means that for any nonzero scalar $c\in\mathbb{R}$, multiplying one layer by $c$ and the next layer by $1/c$, or vice versa, yields the same end-to-end function. This phenomenon forms a manifold for global minimizers in the loss landscape (Dinh et al., 2017). Furthermore, we know that $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ is rank-deficient by at least the order of 11/L1-1/L ; that is, at least 11/L1-1/L of the eigenvalues values of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ are zero (Mulayoff & Michaeli, 2020). This means that the ratio of the manifold dimension to the ambient space dimension increases as $L$ grows.
Projection onto the Hessian Eigenvectors. If we use random projections in visualizations, plots might not be informative to track the optimization dynamics of GD due to the phenomenon caused by the scale invariance. To make contour maps as informative as possible, we choose $\bm{\zeta}$ and $\bm{\gamma}$ to be $\bm{v}_{1}$ and $\bm{v}_{N}$, respectively — the eigenvectors of the largest and smallest eigenvalues of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$.

### v5

@@ Appendix C Proofs from Section 3
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ $=\max_{\begin{subarray}{c}\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}$ (73)
$=\max_{\begin{subarray}{c}\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\operatorname{vec}\Biggl(\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\Biggr]\Biggr)}_{2}^{2}$ (74)
$=\max_{\begin{subarray}{c}\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\operatorname{vec}\Biggl(\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\Biggr)}_{2}^{2}$ (75)
$=\max_{\begin{subarray}{c}\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{u}_{i}}_{2}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{\top}\otimes\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\Biggr]\mathbf{u}_{i}}_{2}^{2}.$ (76)
@@
$\mathbf{K}$ $=\begin{bmatrix}\mathbf{I}\otimes\prod_{j=2}^{L}\mathbf{W}^{*}_{j}\big|{\mathbf{W}_{1}^{*}}^{\top}\otimes\Bigl(\prod_{j=3}^{L}\mathbf{W}^{*}_{j}\Bigr)\big|&\dots&\big|\Bigl(\prod_{j=1}^{L-1}\mathbf{W}_{j}^{*}\Bigr)^{\top}\otimes\mathbf{I}\\ \end{bmatrix}$, (77)
$\mathbf{u}$ $=[\mathbf{u}_{1}^{\top}\quad\mathbf{u}_{2}^{\top}\quad\cdots\quad\mathbf{u}_{L}^{\top}]^{\top}.$ (78)
Then
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ $=\max_{\mathbf{u}:\norm{\mathbf{u}}_{2}=1}2\norm{\mathbf{K}\mathbf{u}}_{2}^{2}$ (79)
$=\sigma_{\max}(\mathbf{K}^{\top}\mathbf{K}).$ (80)
Note that $\sigma_{\max}(\mathbf{K}^{\top}\mathbf{K})=\sigma_{\max}(\mathbf{K}\mathbf{K}^{\top})$, and for any two block matrices $\mathbf{A}$ and $\mathbf{B}$ such that
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}\quad\mathbf{A}_{2}\quad\dots\quad\mathbf{A}_{L}\end{bmatrix}\in\mathbb{R}^{M_{1}\times d},\quad\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}\\ \vdots\\ \mathbf{B}_{L}\\ \end{bmatrix}\in\mathbb{R}^{d\times M_{2}}$ (81)
$\mathbf{AB}=\sum_{i=1}^{L}\mathbf{A}_{i}\mathbf{B}_{i},\quad\mathbf{AB}\in\mathbb{R}^{M_{1}\times M_{2}}.$ (82)
Furthermore, for any matrices $\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D}$ such that the matrix products $\mathbf{AB}$ and $\mathbf{CD}$ are well defined, we have
$(\mathbf{A}\otimes\mathbf{C})(\mathbf{B}\otimes\mathbf{D})=\mathbf{AB}\otimes\mathbf{CD}.$ (83)
Using the fact that $(\mathbf{A}\otimes\mathbf{C})^{\top}=\mathbf{A}^{\top}\otimes\mathbf{C}^{\top}$ together with the previous property, it follows that
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\mathbf{w}^{*})\bigr)=2\sigma_{\max}\Biggl(\sum_{i=1}^{L}\mathbf{B}_{i}^{\top}\mathbf{B}_{i}\otimes\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\Biggr),$ (84)
where $\mathbf{A}_{k}=\prod_{i=k+1}^{L}\mathbf{W}_{i}^{*}$ and $\mathbf{B}_{k}=\prod_{i=1}^{k-1}\mathbf{W}_{i}^{*}$.
@@ C.2 Proof of Corollary 28
According to (17), for any $(\mathbf{L}^{*},\mathbf{R}^{*})\in\Omega$,
$\nabla^{2}\mathcal{L}(\mathbf{L}^{*},\mathbf{R}^{*})[\mathbf{U},\mathbf{V}]=2\norm{\mathbf{L}^{*}\mathbf{U}^{\top}+\mathbf{V}\mathbf{R}^{*\top}}_{F}^{2}.$ (85)
Then
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{L}^{*},\mathbf{R}^{*}))$ $=\max_{\begin{subarray}{c}\mathbf{U},\mathbf{V}\\ \|\mathbf{U}\|_{F}^{2}+\|\mathbf{V}\|_{F}^{2}=1\end{subarray}}2\norm{\mathbf{L}^{*}\mathbf{U}^{\top}+\mathbf{V}\mathbf{R}^{*\top}}_{F}^{2}$ (86)
$\leq\max_{\begin{subarray}{c}\mathbf{U},\mathbf{V}\\ \|\mathbf{U}\|_{F}^{2}+\|\mathbf{V}\|_{F}^{2}=1\end{subarray}}2\bigl(\norm{\mathbf{L}^{*}\mathbf{U}^{\top}}_{F}+\norm{\mathbf{V}\mathbf{R}^{*\top}}_{F}\bigr)^{2}$ (87)
$=\max_{\begin{subarray}{c}\mathbf{u},\mathbf{v}\\ \|\mathbf{u}\|_{2}^{2}+\|\mathbf{v}\|_{2}^{2}=1\end{subarray}}2\Bigl(\|(\mathbf{I}\otimes\mathbf{L}^{*})\mathbf{u}\|_{2}+\|(\mathbf{R}^{*}\otimes\mathbf{I})\mathbf{v}\|_{2}\Bigr)^{2}$ (88)
$\leq\max_{\begin{subarray}{c}\mathbf{u},\mathbf{v}\\ \|\mathbf{u}\|_{2}^{2}+\|\mathbf{v}\|_{2}^{2}=1\end{subarray}}2\Bigl(\sigma_{\max}(\mathbf{I}\otimes\mathbf{L}^{*})\|\mathbf{u}\|_{2}+\sigma_{\max}(\mathbf{R}^{*}\otimes\mathbf{I})\|\mathbf{v}\|_{2}\Bigr)^{2}.$ (89)
We can upper bound the right-hand side of (86) using the triangle inequality. By applying the vectorization trick of the Kronecker product again, we can rewrite (87). Then, noting that for any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ and vector $\mathbf{x}\in\mathbb{R}^{n}$, $\norm{\mathbf{Ax}}_{2}\leq\sigma_{\max}(\mathbf{A})\norm{\mathbf{x}}_{2}$, we can upper bound the right-hand side of (88). Note that for any matrix $\mathbf{A}$ and $\mathbf{B}$, $\sigma_{\max}(\mathbf{A}\otimes\mathbf{B})=\sigma_{\max}(\mathbf{A})\sigma_{\max}(\mathbf{B})$. Hence,
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{L}^{*},\mathbf{R}^{*}))\leq\max_{\begin{subarray}{c}\mathbf{u},\mathbf{v}\\ \|\mathbf{u}\|_{2}^{2}+\|\mathbf{v}\|_{2}^{2}=1\end{subarray}}2\bigl(\sigma_{\max}{(\mathbf{L}^{*})}\norm{\mathbf{u}}_{2}+\sigma_{\max}(\mathbf{R}^{*})\norm{\mathbf{v}}_{2}\bigr)^{2}.$ (90)
@@
$\lambda_{\max}\!\bigl(\nabla^{2}\mathcal{L}(\mathbf{L}^{*},\mathbf{R}^{*})\bigr)\;\leq\;2\bigl(\sigma_{\max}(\mathbf{L}^{*})^{2}+\sigma_{\max}(\mathbf{R}^{*})^{2}\bigr).$ (91)
Now, we will show that this upper bound is achievable. Let us decompose $\mathbf{L}^{*}$ as $\mathbf{U}_{L}\mathbf{\Sigma}_{L}\mathbf{V}_{L}^{\top}$ by SVD, and denote by $\mathbf{u}_{L}$ and $\mathbf{v}_{L}$ the left and right singular vectors corresponding to the largest singular value, respectively. Moreover, decompose $\mathbf{R}^{*}$ as $\mathbf{U}_{R}\mathbf{\Sigma}_{R}\mathbf{V}_{R}^{\top}$ by SVD, and denote by $\mathbf{u}_{R}$ and $\mathbf{v}_{R}$ the left and right singular vectors corresponding to the largest singular value, respectively. We determine a particular $(\mathbf{U}^{*},\mathbf{V}^{*})$ such that it achieves the upper bound while satisfying the constraint $\norm{\mathbf{U}^{*}}_{F}^{2}+\norm{\mathbf{V}^{*}}_{F}^{2}=1$. Choose
$\mathbf{U}^{*\top}=\frac{\sigma_{\max}(\mathbf{L}^{*})}{\sqrt{\sigma_{\max}(\mathbf{L}^{*})^{2}+\sigma_{\max}(\mathbf{R}^{*})^{2}}}\mathbf{v}_{L}\mathbf{u}_{R}^{\top},\quad\mathbf{V}^{*}=\frac{\sigma_{\max}(\mathbf{R}^{*})}{\sqrt{\sigma_{\max}(\mathbf{L}^{*})^{2}+\sigma_{\max}(\mathbf{R}^{*})^{2}}}\mathbf{u}_{L}\mathbf{v}_{R}^{\top}.$ (92)
Using the fact that, for any vectors $\mathbf{x}$ and $\mathbf{y}$ $\norm{\mathbf{x}\mathbf{y}^{\top}}_{F}^{2}=\norm{\mathbf{x}}_{2}^{2}\norm{\mathbf{y}}_{2}^{2}$,
$2\norm{\frac{(\sigma_{\max}(\mathbf{L}^{*})^{2}+\sigma_{\max}(\mathbf{R}^{*})^{2})}{\sqrt{\sigma_{\max}(\mathbf{L}^{*})^{2}+\sigma_{\max}(\mathbf{R}^{*})^{2}}}\mathbf{u}_{L}\mathbf{u}_{R}^{\top}}_{F}^{2}=2(\sigma_{\max}(\mathbf{L}^{*})^{2}+\sigma_{\max}(\mathbf{R}^{*})^{2}).$ (93)
@@ D.1 Visualization of the Contour Map of the Loss Landscape
To study the dynamics of deep matrix factorization, we analyze the trajectories of GD. Previous works have visualized neural network loss landscapes to explore their highly non-convex and non-Euclidean structure (Goodfellow et al., 2014; Li et al., 2018). However, the high dimensionality prevents full visualization. As a result, only 1-D (line) or 2-D (surface) visualizations are available. In this paper, we focus on contour maps of the loss landscape in the vicinity of a global minimum and a methodology employed in prior studies to generate them.
Contour Plots with Random Projections. We want to visualize the loss landscape around a global minimum $\mathbf{w}^{*}\in\mathbb{R}^{N}$. We select two random vectors, $\mathbf{\zeta}$ and $\mathbf{\gamma}$, from $\mathbb{R}^{N}$. Then, for any $K\subset\mathbb{R}^{2}$, we can define the function $p:K\rightarrow\mathbb{R}:$
$p(x,y)=\mathcal{L}(\mathbf{w}^{*}+x\mathbf{\zeta}+y\mathbf{\gamma}),\quad\forall(x,y)\in K,$ (94)
@@
Scale Invariance and Manifolds. Note that our loss function is scale-invariant, which means that for any nonzero scalar $c\in\mathbb{R}$, multiplying one layer by $c$ and the next layer by $1/c$, or vice versa, yields the same end-to-end function. This phenomenon forms a manifold for global minimizers in the loss landscape (Dinh et al., 2017). Furthermore, we know that $\nabla^{2}\mathcal{L}(\mathbf{w}^{*})$ is rank-deficient by at least the order of 11/L1-1/L; that is, at least 11/L1-1/L of the eigenvalues of $\nabla^{2}\mathcal{L}(\mathbf{w}^{*})$ are zero (Mulayoff and Michaeli, 2020). This means that the ratio of the manifold dimension to the ambient space dimension increases as $L$ grows.
Projection onto the Hessian Eigenvectors. If we use random projections in visualizations, plots might not be informative to track the optimization dynamics of GD due to the phenomenon caused by the scale invariance. To make contour maps as informative as possible, we choose $\mathbf{\zeta}$ and $\mathbf{\gamma}$ to be $\mathbf{v}_{1}$ and $\mathbf{v}_{N}$, respectively — the eigenvectors of the largest and smallest eigenvalues of $\nabla^{2}\mathcal{L}(\mathbf{w}^{*})$.